OpenStudy (anonymous):

Show that \[({2 \over (x+1)(x+3)})^2 = {1 \over (x+1)^2} - {1 \over x+1}+ { 1 \over x+3} + {1 \over (x+3)^2}\]

OpenStudy (experimentx):

You will need 4 variables. A/(x+1) + B/(x+1)^2 + C/(x+3) + D/(x+3)^ and the usual process ... and plus it's going to be ugly, best of luck

OpenStudy (experimentx):

*A/(x+1) + B/(x+1)^2 + C/(x+3) + D/(x+3)^2

OpenStudy (anonymous):

Could you help me through this ugly process?

OpenStudy (anonymous):

Please :D

OpenStudy (anonymous):

it's going to be MUCH easier if you solve it on paper... here's quite difficult ! ... you're stuck or you're lazy ? :P

OpenStudy (anonymous):

stuck. not lazy :P

OpenStudy (anonymous):

... Ok, give me a minute :F

OpenStudy (anonymous):

I'll try (if I can) lol

OpenStudy (anonymous):

:D

OpenStudy (experimentx):

yes ... it's quite going to be ugly, ... since it's going be lots of simplification ... @Kreshnik will you help her???

OpenStudy (anonymous):

...I'll try, but as you know , my math sucks LOL

OpenStudy (anonymous):

Thanks for trying :D Your math doesn't suck :p

OpenStudy (experimentx):

keep trying ... this is a 10 or 11 grade question ... it should be pretty easy as you are doing other advanced stuff.

OpenStudy (anonymous):

Haha, it's actually a 12th grade question :/

OpenStudy (experimentx):

anyway keep trying ...it;s just simplification.

OpenStudy (anonymous):

lol, I know, Let me go grab a camera, I'll upload a photo :D

OpenStudy (anonymous):

can you continue? ... (It's not hard,) I have to go, Be right back.

OpenStudy (anonymous):

How did you get the numbers at the top? Wouldn't it be A/... +B/.... +C/.... + D/...?

OpenStudy (experimentx):

@Kreshnik thats not right.

OpenStudy (anonymous):

experiment, how is it done?

OpenStudy (experimentx):

yes, you have to simplify interns of A(x+1)(x+3)^2 +B(x+3)^2 + C(x+1)^2(x+3)+D(x+1)^2 simplify this .... you should practice these kinda of stuff.

OpenStudy (anonymous):

Would you first open the brackets like this?\[({2 \over (x+1)(x+3)})^2---> {4 \over (x+1)(x+3)(x+1)(x+3)}?\]

OpenStudy (anonymous):

ahhhh :@ .. I multiplied, I'll do it again. hold on

OpenStudy (anonymous):

would it be\[{4 \over (x+1)(x+3)(x+1)(x+3)}= {A \over x+1}+{B \over x+3}+{C \over x+1} + {D \over x+3}?\]

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

put square below on B, and C

OpenStudy (anonymous):

How come?

OpenStudy (anonymous):

\[{4 \over (x+1)(x+3)^2(x+1)^2(x+1)}={A \over x+1} +{B \over (x+3)^2}+{C \over (x+1)^2}+ {D \over c+3}?\]

OpenStudy (anonymous):

The last = x+3 sorry*

OpenStudy (anonymous):

Would that be the start? @experimentX

OpenStudy (experimentx):

yes,, best of luck

OpenStudy (anonymous):

well, unfortunately, the math I'm doing, seems to be wrong :( .. don't we just have to solve for x? :(:(

OpenStudy (experimentx):

well, no ... it's collecting coefficients of x's

OpenStudy (anonymous):

but what happens if we solve for x? ... I got 0=0 !!

OpenStudy (anonymous):

\[4=A (x+3)^2(x+1)^2(x+3) + B(x+1)(x+1)^2(x+3)+C(x+1)(x+3)^2(x+3)\]\[+D(x+1)(x+3)^2(x+1)^2\]

OpenStudy (anonymous):

?

OpenStudy (experimentx):

Ah ... no you overdid it.

OpenStudy (anonymous):

So.. How?

OpenStudy (experimentx):

A(x+1)(x+3)^2 +B(x+3)^2 + C(x+1)^2(x+3)+D(x+1)^2 Jeez ... what do kids learn these days at school.

OpenStudy (anonymous):

I'm sorry to take your time, obviously I'm 100% dumb, I wish I could help. :( I'm out .

OpenStudy (anonymous):

Umm. how did you get that?

OpenStudy (experimentx):

@Kreshnik you can give a try ...

OpenStudy (anonymous):

I didn't learn it at school :o...

OpenStudy (experimentx):

Taking LCM

OpenStudy (anonymous):

@experimentX I don't know these kind of tasks, I wish I did. Why should I try, when I'm always wrong :(

OpenStudy (anonymous):

Is there a common lcm for this then?

OpenStudy (anonymous):

Can you explain more how you got that?

OpenStudy (experimentx):

yes (x+1)^2(x+3)^2

OpenStudy (experimentx):